Be sure to answer all parts. The solubility of zinc oxalate is 7.9 x 103 Mat 18°C. Calculate its Ksn x 10 Ksp
Answer #1

At equilibrium:

ZnC2O4 <----> Zn2+ + C2O42-

   s s

Ksp = [Zn2+][C2O42-]

Ksp = (s)*(s)

Ksp = 1(s)^2

Ksp = 1(7.9*10^-3)^2

Ksp = 6.241*10^-5

Answer: 6.2*10^-5